Complete combustion of 6.60 g of a hydrocarbon produced 21.1 g of CO2 and 7.55 g of H2O. What is the empirical formula for the hydrocarbon? Insert subscripts as necessary.

Empirical Formula: CH
1
Expert's answer
2019-09-19T05:11:22-0400

Solution.

First, find the mass of carbon in an unknown hydrocarbon:

"m(C) = \frac{12 \times 21.1}{44} = 5.76g"

Now we find the mass of hydrogen in the unknown hydrocarbon:

"m(H) = \frac{2 \times 7.55}{18} = 0.84g"

Now we find the ratio of the amount of carbon and hydrogen in an unknown substance:

"n(C):n(H) = \frac{5.76}{12} : \frac{0.84}{1} = 0.48:0.84 = 1:1.75"

We found the ratio of carbon and hydrogen atoms, but the values are not true. Therefore, we multiply each number by 8, so that we get integer values:

n(C):n(H) = 8:14

The formula of compound is C8H14.

Answer:

The formula of compound is C8H14.